OpenStudy (anonymous):

What is the refractive index of a solution that has a critical angle of 40.8° with respect to air? The refracting index of air is 1.00.

OpenStudy (anonymous):

 \[\mu1*\sin i = \mu2*\sin r\] \[\mu1 = 1, i = 40.8 \deg, r = 90 \deg, \] hence \[\mu2 = \sin i \] = \[\sin (40.8 * \pi / 180 )\] = 0.6534